The perimeter of a rectangle is 24 inches. The ratio of the length to the width is 2:1. Find the length and the width.
Answer by mananth(16946)

The perimeter of a rectangle is 24 inches. The ratio of the length to the width is 2:1. Find the length and the width.
..
PERIMETER
2*(L+W)= 24
/2
L+W= 12
..
W/L = 2/1
W= 2L
..
substitute w
L+W=12
2L+L=12
3L=12
/3
L=4 the length
..
Width W = 2L = 2*4 = 8
..
m.ananth@hotmail.ca